# Chiral anomaly in the Weyl semimetal TaRhTe4

### v1

@@ II Experimental
The setup for DC electric transport measurements was prepared by gluing the electrodes (silver wires with 50 $\mu m$ thick) to the samples using silver paint (Dupont 4929N) as shown in the inset Figure 2a. All transport measurements were carried out within a temperature range of 8 K to room temperature, using a tested homemade probe and Oxford cryostats equipped with sufficiently large magnets. A Keithley Standard Series 2400 Source Measure Unit (SMU) was used to apply the electric current (field), while a Keithley Nanovoltmeter Model 2182A was utilized to measure the MR signal.
In each experiment, we measured the longitudinal resistivity signal. To investigate the anisotropy of the samples, we studied the total magnetic field dependence up to 10 T for two magnetic field geometries: in-plane (parallel and perpendicular to the electric field) and out-of-plane (perpendicular to the electric field).
@@ Acknowledgements
We would like to thank Tino Schreiner and Danny Baumann for their technical support. MB acknowledges Forschungsstipendien für Doktorandinnen und Doktoranden, 2024 (DAAD) via project number 57694190 and the IFW Excellence Program 2018. SA acknowledges DFG via project number 523/4-1. We would like to acknowledge the financial support provided by the Deutsche Forschungsgemeinschaft (DFG) through Project number 449494427. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Wurzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter—ct.qmat (EXC 2147, project ID 242021). The project was supported by the Leibniz Association through the Leibniz Competition.

### v2

@@ II Experimental
The setup for DC electric transport measurements was prepared by gluing the electrodes (silver wires with 50 $\mu m$ thick) to the samples using silver paint (Dupont 4929N) as shown in the inset Figure 2a. All transport measurements were carried out within a temperature range from 8 K to room temperature, using a tested homemade probe and Oxford cryostats equipped with sufficiently large magnets. A Keithley Standard Series 2400 Source Measure Unit (SMU) was used to apply the electric current, while a Keithley Nanovoltmeter Model 2182A was utilized to measure the voltage signal. A semi-sum of the datapoints acquired with positive and negative electric currents has been systematically performed in all the measurements to eliminate the thermoelectric background.
The magnetoresistance has been measured in magnetic field up to 10 T, applied both in-plane (parallel and perpendicular to the electric field) and out-of-plane (perpendicular to the electric field).
@@ Acknowledgements
We would like to thank Tino Schreiner and Danny Baumann for their technical support. MB acknowledges Forschungsstipendien für Doktorandinnen und Doktoranden, 2024 (DAAD) via project number 57694190 and the IFW Excellence Program 2018. SA acknowledges DFG via project number 523/4-1. We would like to acknowledge the financial support provided by the Deutsche Forschungsgemeinschaft (DFG) through Project number 449494427. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Wurzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter—ct.qmat (EXC 2147, project ID 242021). The project was supported by the Leibniz Association through the Leibniz Competition. Data Availability
The data that support the findings of this article are openly available DataAvailability